OpenStudy (anonymous):

divide. simplify by taking roots if possible sqrt15a/sqrt3a

OpenStudy (unklerhaukus):

\[ {\sqrt{15a} \over \sqrt{3a}}=\sqrt{15a/3a}=√5\]